OpenStudy (anonymous):

Solve: w^2/3=2w^1/3+8 PLEASE HELP ME GUYS! REALLY NEED YOUR HELP!

OpenStudy (anonymous):

\[w^{\frac{2}{3}}-2w^{\frac{1}{3}}-8=0\] is a start

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

SO I WILL ADD 8 TO THE OTHER SIDE

OpenStudy (anonymous):

i meant to each side

OpenStudy (anonymous):

then this is a quadratic equation in \(w^{\frac{1}{2}}\)

OpenStudy (anonymous):

you can do that if you like, and solve by completing the square or you can factor

OpenStudy (anonymous):

there is simpler way let w^1/3=x so equation becomes x^2-2x-8=0 slove for x first

OpenStudy (anonymous):

\[(u-4)(u+2)=0\] where \(u=w^{\frac{1}{2}}\)

OpenStudy (anonymous):

same thing @satellite73 is saying

OpenStudy (anonymous):

then for w

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it x+4=0 and x+2=0?

OpenStudy (anonymous):

no it is \(x-4=0\) or \(x+2=0\)

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

so what will my answer be? my final? (x-4)(x+2)?

OpenStudy (anonymous):

=0?

OpenStudy (anonymous):

no that is just a step along the way your job is to solve for \(w\)

OpenStudy (anonymous):

x=4, x=-2?

OpenStudy (anonymous):

yes and now replace \(x\) by \(\sqrt{w}\)

OpenStudy (anonymous):

i am just messed up

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

both didn't work, i tried

OpenStudy (anonymous):

lets go slow don't forget what you started with: solve for \(w\) if \[w^{\frac{2}{3}}-2w^{\frac{1}{3}}-8=0\]

OpenStudy (anonymous):

we replace \(w^{\frac{1}{3}}\) by \(x\) and get \(x=4\) or \(x=-2\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then go back and say \[w^{\frac{1}{3}}=4\] this means the cube root of \(w\) is 4

OpenStudy (anonymous):

if the cube root of \(w\) is 4, then \(w=4^3=64\)

OpenStudy (anonymous):

and if the cube root of \(w\) is \(-2\) then \(w=(-2)^3=-8\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh i see i made a mistake above in my third post, that was my faulty sorry to confuse

OpenStudy (anonymous):

in any case we get two possible answers for \(w\) either \(w=64\) or \(w=-8\)

OpenStudy (anonymous):

i will try that

OpenStudy (anonymous):

ahh on line class!

OpenStudy (anonymous):

;D

OpenStudy (anonymous):

THE ANSWER IS -8, BECAUSE 64 DIDN'T WORK

OpenStudy (anonymous):

well that is simply wrong 64 works for sure

OpenStudy (anonymous):

HOW COME?

OpenStudy (anonymous):

you can check in the original equation if you put \(w=64\) you get \(4^2=2\times 4+8\) i.e. \(16=8+8\) so 64 is a solution whether or not the computer agrees

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

did you try both?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

YES IT DID WORK. 64 IS THE ANSWER

OpenStudy (anonymous):

I HAVE ON MORE, I DID IT. CAN YOU PLEASE CHECK IT FOR ME?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

WRITE THE EQUATION OF A QUADRATIC WITH SOLUTIONS OF 1/3 AND -1/4. MY ANSWER IS 12 X^2-X-1

OpenStudy (anonymous):

ok lets try it \[(3x-1)(4x+1)=x^2-4x+3x-1=x^2-x-1\] yes you are right

OpenStudy (anonymous):

i mean \[12x^2-x-1\] yes

OpenStudy (anonymous):

your answer is correct good work

OpenStudy (anonymous):

gotta run, good luck